Tohban Report 2011-03-16

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 09 Mar 2011
End Date: 16 Mar 2011
Tohban: Lindsay Glesener and Nicole Duncan
Tohban email: and nicoleduncan@ssl.berkeley.eduwarning.pngURIs of the form " and" are not allowed.
Next Tohban: Hazel Bain
List all reports


Solar Activity

Solar activity is decreased from last week, baseline level is mid-to-high B class, versus C class last week. There was one X-class flare (caught by RHESSI), 4 M-class flares, and several C-class flares. (More specific update once a bug with hsi_tohban_flare_counter is worked out.) The most active regions, 11166 and 11169 will soon rotate off the disk, and the 2-3 ARs remaining have not been emitting flares, and STEREO-B shows that we will not have new ARs on the disk for a few days. This week is therefore expected to be quieter than last.

Memory Management

SSR hit a max of 80% due to the high activity a week ago. The SSR has emptied after the most recent pass; the decrease was due to added passes and the switch to Active/Vigorous decimation made on the 9th.

Data Gaps

There were no data gaps, although the time period ~10 to 11 on March 9 has data that appears to be jumbled due to bad packets.

RHESSI Operations

Commanding was done a few hours after last week's meeting (March 9): Detector 9 HV was lowered by 30 steps, and decimation was changed to ACTIVE / VIGOROUS. Detector 9's LLD (rear) was changed to FF.

Detector 9 continues to show odd behavior, but in a different way from last week. (The change in behavior roughly corresponded to the commanding of March 9.) Instead of large, frequent spikes, there are now few-minute periods of elevated rates in the front segment and decreased rates in the rear segment. (This could be due to a loss of segmentation.) For now, detector 9 controls will be left the same, but the threshold may be changed again next week.

Personal tools